Okay, so in general g(x)=x^2-2x, and the problem is 2(g(t))
Answer by jim_thompson5910(35256)

g(x)=x^2-2x

g(t)=t^2-2t

2*g(t)=2*(t^2-2t)

2*g(t) = 2t^2-4t